OpenStudy (anonymous):

[10.01] Simplify 4 times f to the sixth power times g squared all over 16 times g cubed.. f squared times g squared all over 8. f to the sixth power over 12 times g. f squared times g squared all over 4. f to the sixth power over 4 times g.

OpenStudy (anonymous):

\[\frac{ 4f^6*g^2 }{ 16g^3 }\] \[\frac{ x^a }{ x^b } = x^{a-b}\] \[\frac{ 4 }{ 16 }f^6 g^{2 - 3} = \frac{ f^6 g^{-1} }{ 4 } = \frac{ f^6 }{ 4g }\]